# Strict quantization for compact pseudo-Kähler manifolds and group actions

### v1

@@ 2.1 Standard notations
The following notations are adopted through this article: $\mathbb{N}$ is the set of natural numbers, $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, $\mathbb{R}$ is the set of real numbers, and we denote by +=${\mathbb{R}}_{+}=\{x\in\mathbb{R};\,x>0\}$ and ¯+=$\overline{\mathbb{R}}_{+}=\{x\in\mathbb{R};\,x\geq 0\}$. Furthermore, we write $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}_{0}$ if $\alpha_{j}\in\mathbb{N}_{0}$, $j=1,\ldots,n$.
@@
Now, we recall the Schwartz kernel theorem. Let $B$ and $E$ be smooth vector bundles over paracompact orientable smooth manifolds $M$ and $N$, respectively, equipped with smooth densities of integration. If
$A:{C}^{\infty}_{0}(N,\,E)\rightarrow D^{\prime}(M,\,B)$
is continuous, we write $A(x,y)$ to denote the distribution kernel of $A$. The following two statements are equivalent
  1. 1.
@@
  2. 2.
    The distributional kernel $A(x,y)\in{C}^{\infty}(M\times N,B\boxtimes E^{*})$.
If $A$ satisfies (1) or (2), we say that $A$ is smoothing on $M\times N$. Let
$A,\hat{A}:{C}^{\infty}_{0}(N,\,E)\rightarrow{D}^{\prime}(M,\,B)$
be continuous operators, then we write
$A\equiv\hat{A}$ (on $M\times N$)
if $A-\hat{A}$ is a smoothing operator. If $M=N$, we simply write “on $M$”. We say that $A$ is properly supported if the restrictions of the two projections $(x,y)\mapsto x$, $(x,y)\mapsto y$ to ${\rm supp\,}(A(x,y))$ are proper.
Eventually, let
$H(x,y)\in D^{\prime}(M\times N,\,B\boxtimes E^{*})\,$
then we write $H$ to denote the unique continuous operator
$C^{\infty}_{0}(N,E)\rightarrow D^{\prime}(M,B)$
with distribution kernel $H(x,y)$ and we identify $H$ with $H(x,y)$.
@@ 2.2 Operators and symbols
Let $(X,T^{1,0}X)$ be a compact, connected and orientable CR manifold of dimension $12n+$ 1, $n\geq 1$, we adopt the same notation as in Introduction.
Let $D$ be an open set of $X$. Recall that $\Omega^{0,q}(D)$ denote the space of smooth sections of $T^{*0,q}X$ over $D$ and let $\Omega^{0,q}_{c}(D)$ be the subspace of $\Omega^{0,q}(D)$ whose elements have compact support in $D$. Let
¯$\overline{\partial}_{b}:\Omega^{0,q}(X)\rightarrow\Omega^{0,q+1}(X)$
@@
[…]
¯$\overline{\partial}_{b}:{\rm Dom\,}\overline{\partial}_{b}\subset L^{2}_{(0,r)}(X)\rightarrow L^{2}_{(0,r+1)}(X)\,,$
@@
${\rm Dom\,}\overline{\partial}_{b}:=\{u\in L^{2}_{(0,r)}(X);\,\overline{\partial}_{b}u\in L^{2}_{(0,r+1)}(X)\}$
@@
¯$\overline{\partial}^{*}_{b}:{\rm Dom\,}\overline{\partial}^{*}_{b}\subset L^{2}_{(0,r+1)}(X)\rightarrow L^{2}_{(0,r)}(X)$
[…]
$\Big{\{}s\in L^{2}_{(0,q)}(X);\,s\in{\rm Dom\,}\overline{\partial}_{b}\cap{\rm Dom\,}\overline{\partial}^{*}_{b},\,\overline{\partial}_{b}s\in{\rm Dom\,}\overline{\partial}^{*}_{b},\,\overline{\partial}^{*}_{b}s\in{\rm Dom\,}\overline{\partial}_{b}\Big{\}}$
@@
$\Box^{q}_{b}s=\overline{\partial}_{b}\overline{\partial}^{*}_{b}s+\overline{\partial}^{*}_{b}\overline{\partial}_{b}s\>\>\text{for $s\in{\rm Dom\,}\Box^{q}_{b}$}\,.$
Let
$S^{(q)}:L^{2}_{(0,q)}(X)\rightarrow{\rm Ker\,}\Box^{q}_{b}$ (7)
[…]
$S^{(q)}(x,y)\in{D}^{\prime}(X\times X,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
denote the distribution kernel of $S^{(q)}$.
Now, we recall Hörmander symbol space. Let $D\subset X$ be a local coordinate patch with local coordinates $x=(x_{1},\ldots,x_{2n+1})$.
@@
For $m\in\mathbb{R}$,
$S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
is the space of all
$a(x,y,t)\in{C}^{\infty}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
$\left|\partial^{\alpha}_{x}\partial^{\beta}_{y}\partial^{\gamma}_{t}a(x,y,t)\right|\leq C_{\alpha,\beta,\gamma}(1+\left|t\right|)^{m-\gamma},\ \ \mbox{forevery $(x,y,t)\in K\times\mathbb{R}_{+},t\geq 1$}.$
Put
$S^{-\infty}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
to be
$\bigcap_{m\in\mathbb{R}}S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}).$
Now, let
$a_{j}\in S^{m_{j}}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,,$
$j=0,1,2,\ldots$ with $m_{j}\rightarrow-\infty$, as $j\rightarrow\infty$. Then there exists
$a\in S^{m_{0}}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
unique modulo $S^{-\infty}$, such that
$a-\sum^{k-1}_{j=0}a_{j}\in S^{m_{k}}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}\big{)}$
for $k=0,1,2,\ldots$.
$S^{m_{0}}_{1,0}\big{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}\big{)}\,.$
Furthermore, we write
$s(x,y,t)\in S^{m}_{{\rm cl\,}}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
if
$s(x,y,t)\in S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
and
$s(x,y,t)\sim\sum^{\infty}_{j=0}s^{j}(x,y)t^{m-j}\text{ in }S^{m}_{1,0}(D\times D\times\mathbb{R}_{+}\,,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,,$
where
$s^{j}(x,y)\in{C}^{\infty}(D\times D,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}),\ j\in\mathbb{N}_{0}\,.$
Sometimes we simply write $S^{m}_{1,0}$ to denote
$S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,,$
where $m\in\mathbb{R}\cup\left\{-\infty\right\}$.
Let $D\subset X$ be an open set. Let
$L^{m}_{{\rm cl\,}}(D,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
denote the space of classical pseudodifferential operators on $D$ of order $m$ from sections of $T^{*0,q}X$ to sections of $T^{*0,q}X$ respectively. Let
$P\in L^{m}_{{\rm cl\,}}(D,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,$
and we write $\sigma_{P}$ and $\sigma^{0}_{P}$ to denote the full symbol of $P$ and the principal symbol of $P$ respectively.
2.3 Toeplitz operators
In this sections we briefly recall some results from [GH1]. First let us recall the assumptions from Introduction. Here, the Reeb vector field $R$ is transversal to the space $\underline{\mathfrak{g}}$ at every point $p\in\mu^{-1}(0)$,
$e^{i\theta}\cdot g\circ x=g\circ e^{i\theta}\cdot x,\ \ \mbox{for every $x\in X$, $\theta\in[0,2\pi[$, $g\in G$,}$
and
The action of $G$ and $S^{1}$ is free near $\mu^{-1}(0)$.
$T^{G}_{k}[g]:=S^{G}_{k}\circ g\cdot\circ S^{G}_{k}:L^{2}_{(0,n_{-})}(X)\rightarrow L^{2}_{(0,n_{-}),k}(X)^{G}.$ (8)
The main interest for us is the case when $g=\tilde{f}$ and $f$ is in $C^{\infty}(X_{G})^{S^{1}}$. For the following theorem see Theorem $1.8$ in [HH] and [GH1].
Theorem 2.1.
Let $f\in{C}^{\infty}(X)^{G}$ and suppose $q=n_{-}$. Let $D$ be an open set in $X$ such that the intersection $\mu^{-1}(0)\cap D=\emptyset$. Then,
$T^{G}_{k}[f]\equiv O(k^{-\infty})$
on $D$.
Let $p\in\mu^{-1}(0)$ and let $U$ a local neighborhood of $p$. Then, if $q=n_{-}$, for every fixed $y\in U$, we consider $T_{k}^{G}[f](x,y)$ as a $k$-dependent smooth function in $x$. There exist local coordinates $(x_{1},\dots\,x_{2n+1})$ on $U$ such that
$T^{G}_{k}[f](x,\,y)=e^{ik\,\Psi(x,\,y)}\,b(x,\,y,\,k)+O(k^{-\infty})\,$
for every $x,\,y\in U$. For an explicit expression of the phase function $\Psi$ we refer to [GH1], the symbol satisfies
$b(x,\,y,\,m)\in S^{n-d/2}_{\mathrm{loc}}(1,\,U\times U,\,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,$
$b_{0}(x,\,x):=2^{d-1}\frac{1}{V_{{\rm eff\,}}(x)}\pi^{-n-1+\frac{d}{2}}\left|\det R_{x}\right|^{-\frac{1}{2}}\left|\det{L}_{x}\right|\tau_{x,n_{-}}$
where
$V_{{\rm eff\,}}(x):=\int_{O_{x}}\mathrm{dV}_{O_{x}}\,,$
$O_{x}$ is the orbit through $x$, and $R_{x}$ is the linear map
$\begin{array}[]{rll}R_{x}:\underline{\mathfrak{g}}_{x}&\rightarrow&\underline{\mathfrak{g}}_{x},\\ u&\rightarrow&R_{x}u,\ \ \langle\,R_{x}u\,|\,v\,\rangle=\langle\,\mathrm{d}\omega_{0}(x)\,,\,Ju\wedge v\,\rangle.\end{array}$
Now we recall the definition of the operator $\hat{H}$, which appears in the description of the $C_{j}$’s, see [GH1]. It is convent to introduce a class of operator which looks microlocally like the Szegő kernel. Here we recall the definition from [GH1] for $q=n_{-}$ and action free case.
Definition 2.2.
Let $H:\Omega^{0,n_{-}}(X)\rightarrow\Omega^{0,n_{-}}(X)$ be a continuous operator with distribution kernel
$H(x,y)\in{D}^{\prime}(X\times X,T^{*0,n_{-}}X\boxtimes(T^{*0,n_{-}}X)^{*})\,.$
@@ 2.3 Toeplitz operators
$a_{-}\in S^{k-\frac{d}{2}}_{{\rm cl\,}}(D\times D\times\mathbb{R}_{+},T^{*0,n_{-}}X\boxtimes(T^{*0,q}X)^{*})\,,$
@@
Here we adapt the discussion before Theorem 4.9 in [GH1] to $G$-action case. Let
$\hat{R}:=\frac{1}{2}S^{G}(-iR+(-iR)^{*})S^{G}:\Omega^{0,n_{-}}(X)\rightarrow\Omega^{0,n_{-}}(X),$ (9)
[…]
$\begin{split}&S^{G}\hat{H}=\hat{H}=\hat{H}S^{G}\,,\\ &\hat{R}\hat{H}\equiv\hat{H}\hat{R}\equiv S^{G}\,.\end{split}$ (10)
Note that $\hat{H}$ is uniquely determined by (10), up to some smoothing operators. The operator $\hat{H}$ can always be found by setting
$\hat{H}=S^{G}\tilde{H}S^{G}$
for a given pseudodifferential operator $\tilde{H}$. In fact, the first equation in (10) is always satisfied. Furthermore the second equation can be solved by an easy application of the stationary phase formula.
Now, we can recall the $G$-invariant version of Theorem 4.9 in [GH1].
@@
Recall that we work with the assumption that $q=n_{-}$. Let $f,g\in{C}^{\infty}(X)$, then we have
$T^{G}[f]\circ T^{G}[g]-\sum^{N}_{j=0}\hat{H}^{j}\,T^{G}\left[{\hat{C}_{j}(f,g)}\right]\in\Psi_{n-N-1}^{G}(X),$
for every $N\in\mathbb{N}_{0}$, where $\hat{C}_{j}(f,g)\in{C}^{\infty}(X)$, $\hat{C}_{j}$ is a universal bidifferential operator of order $\leq 2j$, $j=0,1,\ldots$, and
$\begin{split}\hat{C}_{0}(f,g)&=f\cdot g,\\ \hat{C}_{1}(f,g)-\hat{C}_{1}(g,f)&=i\,\{f,g\}.\end{split}$
Eventually, we recall the following theorem from [GH1].
@@
$\left\|T^{G}_{f,k}\circ T^{G}_{g,k}-T^{G}_{g,k}\circ T^{G}_{f,k}-\frac{1}{k}\,T^{G}_{i\{f,g\},k}\right\|=O(k^{-2}),$ (11)
@@
$\left\|T^{G}_{f,k}\circ T^{G}_{g,k}-\sum^{N}_{j=0}k^{-j}\,T^{G}_{C_{j}(f,g),k}\right\|=O(k^{-N-1})$ (12)
in $L^{2}$ operator norm, for every $N\in\mathbb{N}$, where $C_{j}(f,g)\in{C}^{\infty}(X)^{G}$, $C_{j}$ is a universal bidifferential operator of order $\leq 2j$, $j=0,1,\ldots$, and
$\begin{split}C_{0}(f,g)&=f\cdot g\,,\\ C_{1}(f,g)-C_{1}(g,f)&=i\,\{f,g\}\,.\end{split}$ (13)

### v2

@@ 3.1 Standard notations
The following notations are adopted through this article: $\mathbb{N}$ is the set of natural numbers, $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, $\mathbb{R}$ is the set of real numbers, and we denote by +=${\mathbb{R}}_{+}=\{x\in\mathbb{R}\mid x>0\}$ and ¯+=$\overline{\mathbb{R}}_{+}=\{x\in\mathbb{R}\mid x\geq 0\}$. Furthermore, we write $\alpha=(\alpha_{1},\dots,\alpha_{n})\in\mathbb{N}^{n}_{0}$ if $\alpha_{j}\in\mathbb{N}_{0}$, $j=1,\dots,n$.
@@
Now, we recall the Schwartz kernel theorem. Let $B$ and $E$ be smooth vector bundles over paracompact orientable smooth manifolds $M$ and $N$, respectively, equipped with smooth densities of integration. If $A\colon{C}^{\infty}_{0}(N,E)\rightarrow D^{\prime}(M,B)$ is continuous, we write $A(x,y)$ to denote the distribution kernel of $A$. The following two statements are equivalent
  1. (1)
@@
  2. (2)
    the distributional kernel $A(x,y)$ lies in ${C}^{\infty}(M\times N,B\boxtimes E^{*})$.
If $A$ is continuous, we say that $A$ is smoothing on $M\times N$. Let
$A,\hat{A}\colon\ {C}^{\infty}_{0}(N,E)\rightarrow{D}^{\prime}(M,B)$
be continuous operators, then we write $A\equiv\hat{A}$ (on $M\times N$) if $A-\hat{A}$ is a smoothing operator. If $M=N$, we simply write “on $M$”. We say that $A$ is properly supported if the restrictions of the two projections $(x,y)\mapsto x$, $(x,y)\mapsto y$ to ${\rm supp}(A(x,y))$ are proper.
Eventually, let $H(x,y)\in D^{\prime}(M\times N,B\boxtimes E^{*})$ then we write $H$ to denote the unique continuous operator $C^{\infty}_{0}(N,E)\rightarrow D^{\prime}(M,B)$ with distribution kernel $H(x,y)$ and we identify $H$ with $H(x,y)$.
@@ 3.2 Operators and symbols
Let $\bigl{(}X,T^{1,0}X\bigr{)}$ be a compact, connected and orientable CR manifold of dimension $12n+$ 1, $n\geq 1$.
@@
$(u|v):=\int_{X}\langle u(x)|v(x)\rangle{\rm d}v(x),\qquad u,v\in\Omega^{0,q}(X).$
Recall that $L^{2}_{(0,q)}(X)$ denote the completion of $\Omega^{0,q}(X)$ with respect to (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). We extend (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) to $L^{2}_{(0,q)}(X)$ in the standard way. For $f\in L^{2}_{(0,q)}(X)$, we denote $\left\|f\right\|^{2}:=(f|f)$. We extend ¯$\overline{\partial}_{b}$ to $L^{2}_{(0,r)}(X)$, $r=0,1,\dots,n$, by
¯$\overline{\partial}_{b}\colon\ \operatorname{Dom}\overline{\partial}_{b}\subset L^{2}_{(0,r)}(X)\rightarrow L^{2}_{(0,r+1)}(X),$
@@
$\operatorname{Dom}\overline{\partial}_{b}:=\big{\{}u\in L^{2}_{(0,r)}(X)\mid\overline{\partial}_{b}u\in L^{2}_{(0,r+1)}(X)\big{\}}$
@@
¯$\overline{\partial}^{*}_{b}\colon\ \operatorname{Dom}\overline{\partial}^{*}_{b}\subset L^{2}_{(0,r+1)}(X)\rightarrow L^{2}_{(0,r)}(X)$
to denote the Hilbert space adjoint of ¯$\overline{\partial}_{b}$ in the $L^{2}$ space with respect to (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). Let $\Box^{q}_{b}$ denote the Gaffney extension of the Kohn Laplacian whose domain $\operatorname{Dom}\Box^{q}_{b}$ is given by
$\big{\{}s\in L^{2}_{(0,q)}(X)\mid s\in\operatorname{Dom}\overline{\partial}_{b}\cap\operatorname{Dom}\overline{\partial}^{*}_{b},\,\overline{\partial}_{b}s\in\operatorname{Dom}\overline{\partial}^{*}_{b},\,\overline{\partial}^{*}_{b}s\in\operatorname{Dom}\overline{\partial}_{b}\big{\}}$
and $\Box^{q}_{b}s=\overline{\partial}_{b}\overline{\partial}^{*}_{b}s+\overline{\partial}^{*}_{b}\overline{\partial}_{b}s$ for $s\in\operatorname{Dom}\Box^{q}_{b}$. Let
$S^{(q)}\colon\ L^{2}_{(0,q)}(X)\rightarrow\operatorname{Ker}\Box^{q}_{b}$
be the orthogonal projection with respect to the $L^{2}$ inner product (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) and let
$S^{(q)}(x,y)\in{D}^{\prime}\bigl{(}X\times X,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
denote the distribution kernel of $S^{(q)}$.
Now, we recall Hörmander symbol space. Let $D\subset X$ be a local coordinate patch with local coordinates $x=(x_{1},\dots,x_{2n+1})$.
Definition 3.1.
For $m\in\mathbb{R}$, $S^{m}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$ is the space of all
$a(x,y,t)\in{C}^{\infty}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
$\big{|}\partial^{\alpha}_{x}\partial^{\beta}_{y}\partial^{\gamma}_{t}a(x,y,t)\big{|}\leq C_{\alpha,\beta,\gamma}(1+\left|t\right|)^{m-\gamma}\qquad\text{for every}\ (x,y,t)\in K\times\mathbb{R}_{+},\ t\geq 1.$
Put $S^{-\infty}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$ to be
$\bigcap_{m\in\mathbb{R}}S^{m}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}.$
Now, let
$a_{j}\in S^{m_{j}}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)},$
$j=0,1,2,\dots$ with $m_{j}\rightarrow-\infty$, as $j\rightarrow\infty$. Then there exists
$a\in S^{m_{0}}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
unique modulo $S^{-\infty}$ such that
$a-\sum^{k-1}_{j=0}a_{j}\in S^{m_{k}}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
for $k=0,1,2,\dots$.
$S^{m_{0}}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}.$
Furthermore, we write
$s(x,y,t)\in S^{m}_{{\rm cl}}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
if
$s(x,y,t)\in S^{m}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
@@
$s(x,y,t)\sim\sum^{\infty}_{j=0}s^{j}(x,y)t^{m-j}\qquad\text{in}\ S^{m}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)},$
where
$s^{j}(x,y)\in{C}^{\infty}\bigl{(}D\times D,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)},\qquad j\in\mathbb{N}_{0}.$
Sometimes, we simply write $S^{m}_{1,0}$ to denote $S^{m}_{1,0}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$, where $m\in\mathbb{R}\cup\left\{-\infty\right\}$.
Let $D\subset X$ be an open set. Let $L^{m}_{{\rm cl}}\bigl{(}D,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$ denote the space of classical pseudodifferential operators on $D$ of order $m$ from sections of $T^{*0,q}X$ to sections of $T^{*0,q}X$ respectively. Let $P\in L^{m}_{{\rm cl}}\bigl{(}D,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$ and we write $\sigma_{P}$ and $\sigma^{0}_{P}$ to denote the full symbol of $P$ and the principal symbol of $P$, respectively.
3.3 Toeplitz operators
In this subsection, we briefly recall some results from [7]. Here, the Reeb vector field $R$ is transversal to the space $\underline{\mathfrak{g}}$ at every point $p\in\mu^{-1}(0)$, ${\rm e}^{{\rm i}\theta}\cdot g\circ x=g\circ{\rm e}^{{\rm i}\theta}\cdot x$, for every $x\in X$, $\theta\in[0,2\pi]$, $g\in G$, and the action of $G$ and $S^{1}$ is free near $\mu^{-1}(0)$.
$T^{G}_{k}[g]:=S^{G}_{k}\circ g\cdot\circ S^{G}_{k}\colon\ L^{2}_{(0,n_{-})}(X)\rightarrow L^{2}_{(0,n_{-}),k}(X)^{G}.$
The main interest for us is the case when $g=\tilde{f}$ and $f$ is in $C^{\infty}(X_{G})^{S^{1}}$. For the following theorem see in [11, Theorem 1.8] and [7].
Theorem 3.2.
Let $f\in{C}^{\infty}(X)^{G}$ and suppose $q=n_{-}$. Let $D$ be an open set in $X$ such that the intersection $\mu^{-1}(0)\cap D=\varnothing$. Then, $T^{G}_{k}[f]\equiv O(k^{-\infty})$ on $D$.
Let $p\in\mu^{-1}(0)$ and let $U$ a local neighborhood of $p$. Then, if $q=n_{-}$, for every fixed ${y\in U}$, we consider $T_{k}^{G}[f](x,y)$ as a $k$-dependent smooth function in $x$. There exist local coordinates $(x_{1},\dots,x_{2n+1})$ on $U$ such that
$T^{G}_{k}[f](x,y)={\rm e}^{{\rm i}k\Psi(x,y)}b(x,y,k)+O(k^{-\infty})$
for every $x,y\in U$. For an explicit expression of the phase function $\Psi$ we refer to [7], the symbol satisfies
$b(x,y,m)\in S^{n-d/2}_{\mathrm{loc}}\bigl{(}1,U\times U,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
$b_{0}(x,x):=2^{d-1}\frac{1}{V_{{\rm eff}}(x)}\pi^{-n-1+\frac{d}{2}}\left|\det R_{x}\right|^{-\frac{1}{2}}\left|\det{L}_{x}\right|\tau_{x,n_{-}},$
where $V_{{\rm eff}}(x):=\int_{O_{x}}{\rm dV}_{O_{x}}$, $O_{x}$ is the orbit through $x$, and $R_{x}$ is the linear map
$R_{x}\colon\ \underline{\mathfrak{g}}_{x}\rightarrow\underline{\mathfrak{g}}_{x},\qquad u\rightarrow R_{x}u,\qquad\langle R_{x}u|v\rangle=\langle{\rm d}\omega_{0}(x),Ju\wedge v\rangle.$
Now we recall the definition of the operator $\hat{H}$, which appears in the description of the $C_{j}$’s, see [7]. It is convenient to introduce a class of operator which looks microlocally like the Szegő kernel. Here we recall the definition from [7] for $q=n_{-}$ and action free case.
@@ 3.3 Toeplitz operators
Let $H\colon\Omega^{0,n_{-}}(X)\rightarrow\Omega^{0,n_{-}}(X)$ be a continuous operator with distribution kernel
$H(x,y)\in{D}^{\prime}\bigl{(}X\times X,T^{*0,n_{-}}X\boxtimes\bigl{(}T^{*0,n_{-}}X\bigr{)}^{*}\bigr{)}.$
@@
$a_{-}\in S^{k-\frac{d}{2}}_{{\rm cl}}\bigl{(}D\times D\times\mathbb{R}_{+},T^{*0,n_{-}}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)},$
@@
Here we adapt the discussion before [7, Theorem 4.9] to $G$-action case. Let
$\hat{R}:=\frac{1}{2}S^{G}(-{\rm i}R+(-{\rm i}R)^{*})S^{G}\colon\ \Omega^{0,n_{-}}(X)\rightarrow\Omega^{0,n_{-}}(X),$
where $(-{\rm i}R)^{*}$ is the adjoint of $-iR$ with respect to (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). Then, we note that $\hat{R}\in\Psi_{n+1}(X)^{G}$ with $\sigma^{0}_{\hat{R},-}(x,x)\neq 0$, for every $x\in X$. Let $\hat{H}\in\Psi_{n-1}(X)^{G}$ with
$S^{G}\hat{H}=\hat{H}=\hat{H}S^{G},\qquad\hat{R}\hat{H}\equiv\hat{H}\hat{R}\equiv S^{G}.$ (3.1)
Note that $\hat{H}$ is uniquely determined by (3.1), up to some smoothing operators. The operator $\hat{H}$ can always be found by setting $\hat{H}=S^{G}\tilde{H}S^{G}$ for a given pseudodifferential operator $\tilde{H}$. In fact, the first equation in (3.1) is always satisfied. Furthermore, the second equation can be solved by an easy application of the stationary phase formula.
Now, we recall the $G$-invariant version of [7, Theorem 4.9].
@@
Let $q=n_{-}$. Let $f,g\in{C}^{\infty}(X)$, then we have
$T^{G}[f]\circ T^{G}[g]-\sum^{N}_{j=0}\hat{H}^{j}T^{G}\big{[}{\hat{C}_{j}(f,g)}\big{]}\in\Psi_{n-N-1}^{G}(X)$
for every $N\in\mathbb{N}_{0}$, where $\hat{C}_{j}(f,g)\in{C}^{\infty}(X)$, $\hat{C}_{j}$ is a universal bidifferential operator of order $\leq 2j$, $j=0,1,\dots$, and $\hat{C}_{0}(f,g)=f\cdot g$, $\hat{C}_{1}(f,g)-\hat{C}_{1}(g,f)={\rm i}\{f,g\}$.
Eventually, we recall the following theorem from [7].
@@
$\left\|T^{G}_{f,k}\circ T^{G}_{g,k}-T^{G}_{g,k}\circ T^{G}_{f,k}-\frac{1}{k}T^{G}_{{\rm i}\{f,g\},k}\right\|=O\bigl{(}k^{-2}\bigr{)},$
@@
$\left\|T^{G}_{f,k}\circ T^{G}_{g,k}-\sum^{N}_{j=0}k^{-j}T^{G}_{C_{j}(f,g),k}\right\|=O\bigl{(}k^{-N-1}\bigr{)}$
in $L^{2}$ operator norm, for every $N\in\mathbb{N}$, where $C_{j}(f,g)\in{C}^{\infty}(X)^{G}$, $C_{j}$ is a universal bidifferential operator of order $\leq 2j$, $j=0,1,\dots$, and $C_{0}(f,g)=f\cdot g$, $C_{1}(f,g)-C_{1}(g,f)={\rm i}\{f,g\}$.